### v1

# Amorphous silicon structures generated using a moment tensor potential and the activation relaxation technique nouveau

Abstract
Abstract
Preparing realistic atom-scale models of amorphous silicon (a-Si) is a decades-old condensed matter physics challenge. Herein, we combine the Activation Relaxation Technique nouveau (ARTn) to a Moment Tensor Potential (MTP) to generate seven a-Si models containing between 216 and 4096 atoms. A thorough analysis of their short-range and medium-range structural properties is performed, alongside assessments of excess energy and mechanical properties. The seven ARTn-MTP models are compared with available experimental data and other high quality a-Si models present in the literature. The seven ARTn-MTP a-Si models are in excellent agreement with available experimental data. Notably, several of our models, including the 216-atom, 512-atom, and 1000-atom a-Si models, exhibit low coordination defects without any traces of crystalline grains. Historically overlooked in previous research, our study underlines the need to assess the validity of the continuous random-network hypothesis for the description of perfect amorphous model by characterizing local crystalline environment and to explore the crystallisation process of a-Si through modelling.
I Introduction
Amorphous silicon (a-Si) is used in applications that include photovoltaics, thin-film transistors, battery electrodes, and liquid-crystal displays 1, 2, 3, 4, 5. Recent advances include its use as coating materials for next-generation gravitational wave detectors 6. Additionally, a-Si serves as the canonical disordered solid prototype for the development of modeling methods.
a-Si 7, 8 is idealized as a continuous random network (CRN) 9, characterized by a local atomic environment akin to crystalline silicon 10, 11. It maintains a local tetrahedral crystalline environment without long-range order. Real a-Si, however, is believed to contain defects associated with its disordered nature at a level that depends on its preparation12, 13, 14, 15, 6. These include local coordination defects, medium-range ring defects and two-level systems–the latter contribute to noise in sensitive gravitational wave detectors and cause decoherence in quantum computers 15, 6. Variations in tetrahedral geometry and bond lengths also significantly affect a-Si’s performance. Controlling structural anomalies during preparation poses challenges from both an experimental and a theoretical perspective. Indeed, the structure of a-Si is highly influenced by preparation methods 16; for instance, chemical vapor deposition 17 and ion implantation 18, 19, 13 can introduce voids and coordination defects. Thermal annealing 18, 20, 13, 21 often refines these imperfections, enhancing material quality.
Extensive theoretical research has focused on developing realistic structural models for a-Si 10, 22, 23, 24, 25, 26, 27. Various methods for simulating materials, including a-Si, can be categorized into dynamic and non-dynamic approaches 7, 8. Dynamic methods, traditionally based on density functional theory (DFT) 28 or semiempirical force fields 29, primarily use melt-quench processes. However, these methods face challenges: experimental a-Si cannot be produced by melt-and-quench, DFT simulations are limited in system size and quenching rate, and semi-empirical force fields are often compromised by limited accuracy and transferability issues.
Non-dynamic methods include the Reverse Monte Carlo (RMC) approach 30, 31, 32, 33, 34. In RMC, the computer-generated a-Si structure is randomly modified in order for its simulated structure factor and radial distribution function g(r) to match experimental data. RMC’s main drawback is that it can yield multiple configurations for the same g(r) — in other word the solution is degenerate. To overcome this, hybrid RMC methods incorporate force field interactions, ensuring generated configurations are both consistent with experimental data and physically realistic. Variants like the force-enhanced atomic refinement 35, 36 and experimentally constrained molecular relaxation 37 utilize semi-empirical and DFT interactions, respectively. Other key non-dynamic methods include the Wooten-Winer-Weaire (WWW) bond-switching algorithm 10, which optimizes atomic configurations by minimizing energy, and the Activation-Relaxation Technique nouveau (ARTn) 22, 38, 39, 40, which combines activation and relaxation steps to discover low-energy structures. ARTn can be paired with various force fields and ab initio methods, making it suitable for applications that include relaxation of disordered systems and diffusion processes 41, 15, 42, 43, 44, 45, 46, 47, 48, 49.
One of the key components in both dynamical and non-dynamical methods is the interaction potential. The ongoing challenge of balancing accuracy and computational cost in these methods has led to the development of machine learning interatomic potentials (MLIPs) 50, 51, 52. MLIPs combine descriptors of the atomic environment, regression methods, and DFT data. Many MLIPs frameworks have been developed, including artificial neural networks 50, 53, kernel-based methods 54, 55, and linear regression approaches 56, 57, among others. For a detailed description of MLIPs, including their implementation and testing, the reader is referred to the existing literature 52. In the case of a-Si, structures have been modeled using artificial neural networks 58, 59, 60 and Gaussian Approximation Potentials (GAP) 54, 55. In initial studies, the GAP potential was applied to a-Si systems containing between 512 and 4,096 atoms 24. Later, the GAP model was extended to larger systems of up to 100,000 atoms to investigate the the atomistic mechanisms underlying various structural transitions in disordered silicon 61. Other MLIPs, such as the Moment Tensor Potential 57 (MTP), have also been applied to a-Si 62. In these studies, melt-quench MD simulations were employed to generate a-Si structures.
A natural next step is to couple ARTn 22, 38, 39, 40 with machine learning potential to generate a-Si models. The objective of this paper is to generate a-Si models using ARTn-ML, which could enhance our understanding of the a-Si model and open new avenues for further research. The manuscript is organized into three main sections. The Methods section outlines the principles of the ARTn method and details our input models, the ARTn-MTP simulation process, and the analytical tools used. The Results and Discussion Section presents findings from our analysis of the configurations generated by ARTn-MTP. Finally, the key insights and discoveries of our study are summarized in the conclusions.
II Methods
II.1 Activation Relaxation Technique nouveau
ARTn 22, 38 is a computational method used in the study of activated processes, particularly in the context of diffusion and transitions between states in condensed matter systems. ARTn is designed to simulate rare events that signify substantial changes in the microscopic structure of atomic systems, based on the efficient exploration of potential energy landscapes. ARTn employs an event-driven approach in which the system evolves by transitioning between different states or configurations. The concept is based on the understanding that within the configurational space, there is a distinct set of points representing saddle points and adjacent minima in the potential energy surface of the system. As the name implies, ARTn can be summarized as a two-step process 22: activation and relaxation. In the activation step, starting from a local energy minimum, the configuration is pushed toward a nearby first-order saddle point. In the relaxation step, the system is driven over the saddle point and then relaxed to a new energy minimum. The new configuration is accepted based on the Metropolis acceptance probability, which is given by $P_{\text{accept}}=\min\left(1,\exp\left(-\frac{\Delta E}{kT}\right)\right)$, where T is the fictitious temperature, set to 0.25 eV for our simulation, k is the Boltzmann constant, and $\Delta E$ is the energy difference between the old and new configurations 22, 63, 23. The search is restricted to paths that include only saddle points. The ARTn method has demonstrated its value in revealing rare events that conventional simulation methods may struggle to capture due to the slow dynamics and relaxation inherent in these systems. Detailed descriptions of the method can be found elsewhere 22, 38, 39, 64, 65.
II.2 A moment tensor potential for silicon
In practical terms, ARTn serves as a saddle-point search engine, employing an interaction potential as both a force and energy calculator. The simulation’s accuracy is directly influenced by the selection of the force field. Previous studies have relied on semi-empirical force field 42, 22, 66, 15, 67 and ab initio methods 68, 69, 44. Here we employ a machine learning-based forcefield. This potential was developed to jointly describe Si, O, and SiO2 systems 62, and is an improvement over a previously reported Si MTP 70. This potential is based on the MTP framework 57, 71. A detailed description of the development and implementation of this MTP can be found in our previous work 62, 70.
Models name Input configurations for a-Si preparation N atoms a-Si preparation methods
216-R1 Randomly filled 216 ARTn runs using MTP potential
216-R2 Randomly filled 216 ARTn runs using MTP potential
216-FEAR Randomly filled 216 Force hence atomic refinement 36
512-R1 Randomly filled 512 ARTn runs using MTP potential
512-R2 Randomly filled 512 ARTn runs using MTP potential
512-MD Randomly filled 512 MD simulation using MTP potential
512-RMC Randomly filled 512 Reverse monte carlo simulation 72, 24
512-INV Randomly filled 512 Invariant environment refinement technique 72, 24
512-FEAR Randomly filled 512 Force hence atomic refinement 36
512-WWW Randomly filled 512 Bond switching algorithm 72, 24
1000-R Randomly filled 1000 ARTn runs using MTP potential
4096-R-MD Randomly filled + MD runs 4096 ARTn runs using MTP potential
4096-WWW Randomly filled 4096 Bond switching algorithm 36
100k-GAP18 Unknown 100000 MD simulation using GAP potential 61
1000-aSi-SW Amorphous configuration 22, 38, 68 1000 ARTn runs using MTP potential
Tab. 1: A summary of the models discussed in this study. The preparation methods for the input configurations, the number of atoms in each model, and the preparation methods for the final a-Si models are provided. Models without a reference were prepared for this work following the protocol discussed in the methods section.
Fig. 1: Key structural elements characterizing a-Si: (a) bond angle ($\theta$), coordination defects such as dangling bonds (3-fold bonded atoms) and floating bonds (5-fold bonded atoms), and ring structures, including 4-, 5-, 6-, and 7-membered rings. (b) A 100% 4-fold fully coordinated amorphous structure comprising 216 atoms, generated by ARTn-MTP simulation , with 14.40 atoms in a crystalline environment. The atoms shown in yellow within the box are in an amorphous environment, while those colored differently are in a crystalline environment, specifically cubic or hexagonal diamond.
II.3 Protocol for input models preparation and simulation details
We prepared models containing 216, 512, 1000, and 4096 atoms. For the model with 216 atoms, we used two different simulation boxes corresponding to densities of 2.20 g/cm³ (216-R1) and 2.28 g/cm³ (216-R2), respectively. These two boxes were randomly filled with an overlap distance (initial atomic separation) of 2.3 Å for 2.20 g/cm³ and 2 Å for 2.28 g/cm³. For the 512-atom model, we employed two simulation boxes randomly filled with atoms, using the same densities as those applied to the 216-atom model. The first 512-atom model had a density of 2.20 g/cm³ (512-R1) with an initial overlap of 2.3 Å. The second 512-atom model (512-R2) along with the 1000-atom model (1000-R) were randomly filled with a density of 2.28 g/cm³ and an initial overlap of 2 Å. The search for events during ARTn-MTP simulations was initiated using configurations that were already at local minima. Herein, the ARTn simulations begins by applying random displacements to the chosen central atom and its neighbors, based on a user-defined local cut-off radius. Once the displacements are introduced, the algorithm searches for convergence to a saddle point, guided by the lowest curvature. After locating the saddle point, the minimization process begins to find the new minimum. The minimization is carried out using either the Steepest Descent (SD) or the Fast Inertial Relaxation Engine (FIRE) algorithm. In all ARTn-MTP simulations, we used a fictitious temperature of 0.25 eV in the Metropolis accept-reject criterion. Each simulation started with event searches initiated with a local cut-off radius of 3 Å, which the algorithm used to identify the region around the selected central atom which will be deformed. This cut-off radius was progressively increased to a maximum cut-off 5.5 Å in order toexpand the local zone of initial deformaton when the ARTn moves no longer led to energy decrease. The coordination defects level was monitored during the simulations. For coordination defect levels above 5 %, the full list of atoms was used as potential central atoms for ARTn moves. For defect levels at or below 5 %, only atoms with a non-four-fold coordination were allowed to serve as the central atom in ARTn moves.
II.4 Structural analysis
We employ a set of criteria to assess the quality of our amorphous structures. These criteria are derived from considerations of short-range structures, coordination defects, and intermediate-range order. Understanding these different structural aspects is crucial in the study of materials properties, as they influence the physical, chemical, and mechanical behavior of materials. We conduct coordination analysis using OVITO 73 with a bond-length cutoff of 2.85 Å. The analysis of the ring size distribution, providing insights into intermediate-range order, is performed using the R.I.N.G.S code 74 based on the King criteria. We employed the ISAACS code 75 for the calculation of the structure factor while the radial distribution function and angular distribution function were determined using LAMMPS 76 through pseudo simulations.
We also measure the amount of crystallinity in the models. Here, crystallity refers to environments that are cubic or hexagonal diamond to second neighbour, all atoms at the center of this environment or within the first- and second-neighbour shell of a perfectly atom are considered as having crystallinity. We use method developed by Ref. 77 as implemented in OVITO 73.
III Results and discussion
Fig. 2: Coordination numbers of input configurations and ARTn-MTP optimized configurations (final configurations). All non-4-fold coordinated atoms, including under- and over-coordinated atoms, are considered defective. The models analyzed include 216, 512 , 1000 and 4096 atoms.
Tab. 2: Structural characteristics of ARTn-MTP generated amorphous models. This includes density, the positions of the first (r1) and second (r2) peaks of the RDF, average coordination number, average bond angle (with variability indicated in parentheses), and the percentages of 3-fold, 4-fold, and 5-fold coordinated atoms in each mode. A comparison is provided with previously published models. Bold fonts are used to highlight models where fewer than 3 % of atoms were mis-coordinated. Crystallinity is defined as discussed in the Methods section. Cells are left empty when information is not available for models from the literature.
Model density r1 r2 Coordination number $\langle\theta\rangle$ ($\sigma$) 3-fold (%) 5-fold (%) 4-fold (%) Crystallinity (%)
216-R1 2.20 2.36 3.77 3.990 109.20 (9.95) 0.93 0.00 99.07 0.00
216-R2 2.28 2.36 3.70 4.000 109.19 (10.02) 0.00 0.00 100.00 14.40
216-FEAR 36 2.33 2.36 3.81 4.028 108.52 (15.59) 1.39 4.17 94.44
512-R1 2.20 2.36 3.90 4.008 109.11 (11.39) 0.59 1.37 98.04 0.00
512-R2 2.28 2.36 3.90 4.015 108.99 (11.76) 0.19 1.76 98.04 0.00
512-MD 2.28 2.37 3.83 3.996 109.05 (10.95) 1.37 0.98 97.65 3.30
512-RMC 72, 24 2.30 2.34 3.66 4.039 108.57 (14.16) 1.56 5.47 92.96 3.30
512-INV 72, 24 2.31 2.34 3.86 4.054 108.56 (14.47) 1.95 7.42 90.62 0.00
512-WWW 72, 24 2.27 2.34 3.94 4.008 109.05 (10.36) 0.19 0.98 98.82 10.00
512-FEAR 36 2.33 2.35 3.82 4.008 1.17 2.73 95.90
512-GAP18 24 2.27 2.36 3.76 4.004 109.20 (9.60) 0.58 0.98 98.44 6.4
1000-R 2.28 2.37 3.76 4.020 108.94 (12.31) 0.6 2.6 97.20 0.00
4096-R-MD 2.28 2.37 3.83 4.012 109.14 (10.65) 0.51 1.44 98.05 6.40
4096-WWW 36 2.33 2.36 3.78 4.004 0.05 0.49 99.46
100k-GAP18 61 2.27 2.36 3.83 4.003 109.17 (9.99) 0.69 0.95 98.35 6.5
Fig. 3: Comparison of the local structural properties of input configurations (top panels: a, b, and c) with the final configuration (100% coordinated) (bottom panels: d, e, and f) generated using the ARTn-MTP coupling scheme. The structural properties analyzed include angular distribution, radial distribution, and bond length distribution for a system comprising 216 atoms. Experimental data and data from the 100k-GAP18 model 61 are included to assess the quality of our 100% 4-fold coordinated 216-atom a-Si model
Models of size 216 atoms, 512 atoms, 1000 atoms and 4096 atoms were generated as detailed in table 1. We analyze these configurations with respect to short-range order, medium-range order, electronic properties as well as mechanical properties. Fig. 2 displays the defects in both the input and final configurations. The input configurations exhibit a higher prevalence of coordination defects than the final configurations: approximately 70% for the 216-R1, 216-R2, and 512-R1 models, compared to around 30% for the 512-R2 and 1000-R models, as illustrated by the blue bars. The green bars represent the total number of defects present in the final configurations.
In all ARTn-MTP a-Si models, the final percentage of defective atoms remained below 2%, with the exception of the 1000-R and 4096-R-MD models, which exhibit a final defect percentage of approximately 3%. The initial 4-fold coordination number (shown by the orange bars in Fig. 2) is approximately 30% for the 216-R1, 216-R2, and 512-R1 models, and around 68% for the 512-R2 and 1000-R models. In contrast, the final configurations display markedly higher 4-fold coordination numbers (depicted by the red bars in Fig. 2 and Table 2), with values of 99.07%, 100%, 98.04%, 98.04%, 97.20%, and 97.80% for the 216-R1, 216-R2, 512-R1, 512-R2, 1000-R and 4096-R-MD models, respectively. No atoms with fewer than 3 neighbors or more than 5 neighbors were observed.
To further contextualize the number of coordination defects in the ARTn-MTP a-Si models, we present a comparison in Fig. 5, which includes data from the literature as well. First, all the models presented show no atoms with fewer than 3 neighbors or more than 5 neighbors. Secondly, all of the ARTn-MTP generated models exhibit fewer floating and dangling bonds compared to all other methods, except the WWW model and 100k-GAP18 model. Notably, the 216-R2 model is absent from Fig. 5 because it contains 100% four-fold coordination. To our knowledge, no other a-Si models exhibit 100% four-fold coordination, except for some WWW models generated with a Keating-type potential 78, 23
The last column of Table 2 presents the fraction of atoms in a crystalline environment (cubic or hexagonal diamond). As shown, most of the generated a-Si models, particularly those from MD simulations, contain some proportion of crystallinity. Although WWWW is known to produce large a-Si models without coordination defects, the WWW a-Si models still exhibit traces of crystallinity. We identified four models—216-R1, 512-R1, 512-R2, and 1000-R—that are free from crystallinity. Furthermore, as indicated in Table  3, both the 1000-aSi-SW initial and 1000-aSi-SW final configurations also contain no crystallinity. It is important to note that all four of our a-Si models without crystallinity, including the 1000-aSi-SW initial model, were generated starting from a random state without any prior MD prerun, meaning their boxes were initially filled randomly. From Table 2, we note three potential trends. (i) this fraction does not seem to depend much on the system size. (ii) Starting ARTn from a random state seems to lead to a less crystalline state than starting from a melt-and-quench state. (iii) Fewer coordination defects seem to favor a higher crystallinity content. Now that, with reliable ML potentials, we can generate similar quality models using different pathways (non-dynamic melt-and-quench methods), we can start examining the link between local crystallinity, defects and the ideal continuous-random networks. This can lead to understanding why, for example, amorphous silicon cannot be produced experimentally from the liquid phase and crystallizes well below melting as well as whether the CRN model corresponds really to ideal a-Si. The ML potential 62, 70, which leads to high-quality models, makes further study on this question now possible.
Fig. 4: Local structural properties of ARTn-MTP generated a-Si models: (a) radial distribution, (b) angular distribution, and (c) bond length distributions. Also shown are the properties obtained from experimental data and the 100k-GAP18 atom simulations.
Fig. 5: Proportion of coordination defects in various a-Si models generated by different methods. 216-R1 has no floating bond and 216-R2 is 100 % fully 4-fold coordinated. No atoms with fewer than 3 neighbors or more than 5 neighbors have been recorded in any model.
Fig. 6: FWHM of the bond-angle distribution. The FWHM values are computed by fitting a Gaussian distribution to the bond angle distributions. The red line denotes the experimental value of FWHM 20.
III.1 Short-range order structural analysis of the models
Fig. 3 illustrates the structural properties of the 216-R2 models, including the angular distribution function, radial distribution function, and bond length distribution. Panels a), b), and c) display these properties for the randomly packed input configuration (216-R2), while panels d), e), and f) show the properties of the final configurations (216-R2) produced by ARTn-MTP. The final 216-R2 configuration is 100% fully coordinated, with no coordination defects present. As shown in Fig. 3 a), b), and c), the data for the input configurations are randomly distributed and widespread. In contrast, the data for the final configurations, depicted in Fig. 3 d), e), and f), follow well-defined distributions. The shape of the radial distribution function matches the experimental data, though the peak is somewhat overestimated. This indicates that the Si-Si pair distances in our model are widely distributed around the experimental pair separation distance. The bond angle distribution function is centered around the tetrahedral value of 109.5° with a spread of 9.96°. A Gaussian fit to the data reveals a full width at half-maximum (FWHM) of 23.45°. The half width at half-maximum (HWHM) is 11.73°, which is similar to the experimental HWHM value of 11° 20. For the bond length, the data for the input configuration (3 c) are widespread. In contrast, for the final configuration (3 f), the bond lengths follow a Gaussian distribution centered around the experimental value of 2.35 Å. The mean bond length is 2.36 Å, with a spread of 0.03 Å. The angular distribution function, radial distribution function, and bond length distribution of the final model (216-R2) were compared with those of the 100k-GAP18 configuration, which contains 100,000 atoms. As shown in Fig. 3 (d), (e), and (f), these properties align well with those of the larger 100k-GAP18 model. This demonstrates the high quality of our fully coordinated a-Si model. Beyond the analysis of the fully coordinated configuration, we also explore the short-range order structure in our other models. The detailed results are presented in Fig. 4 and Tab. 2. Once again, the structural properties—including the radial distribution function (Fig. 4 a), bond angle distribution function (Fig. 4 b), and bond length distribution (Fig. 4 c)—closely match experimental data and those of 100k-GAP18 models. The data for all the models presented in Fig. 4 show similar behavior, with the exception of the Si-Si bond length distribution. Notably, the bond lengths for two of our models—216-R1 (in green) and 512-R1 (in red)—are shifted. These two models have a density of 2.20 g/cm³, which is lower than that of the other models, which is 2.28 g/cm³. Since amorphous silicon features local tetrahedral bonding, we quantify the deviation from ideal tetrahedral geometry by calculating the FWHM for all models included in this study, as presented in Fig. 6. The FWHM values for all our models closely align with the experimental data, as well as with the WWW model and the 100k-GAP18 models.
III.2 Medium-range order structural analysis of the models
[…]
Fig. 7: Structure factor for ARTn-MTP generated models compared with that obtained from experimental data and 100k-GAP18 atom simulations.
Fig. 8: Ring statistics for ARTn-MTP generated models as well as those derived from the 512-WWW model and 100k-GAP18 atom models.
Ring statistics are illustrated in Fig. 8. Crystalline silicon, characterized by its diamond-type structure, primarily features 6-membered rings, while amorphous silicon tends to favor rings of sizes 5, 6, and 7. Rings with fewer than 5 members or more than 7 are typically regarded as ring defects. Notably, no 3-membered rings are observed in any of the models. Additionally, very few 4-membered rings, which are known to introduce structural stress, were recorded across all models. As shown in Fig. 8 (a-f), each model predominantly contains 5, 6, and 7-membered rings, consistent with existing literature. These ring sizes are energetically favorable for a-Si. Furthermore, all models exhibit a few larger ring defects with sizes greater than 7, as detailed in Fig. 8.
III.3 Refinement of ART-SW optimized model: 1000-aSi-SW
Fig. 9: Local structural properties of ARTn-MTP refined configurations (1000-aSi-SW final) compared to ART-SW optimized configurations (1000-aSi-SW initial). a) Initial and final counts of 3-fold and 5-fold defects, b) bond angle distribution function, c) radial distribution function, and d) structure factor. For figures (b), (c), and (d), experimental data and results from the 100k-GAP-18 atom model are included to assess the quality of both the initial and final configurations of the 1000-aSi-SW model. The ART-SW optimized configurations served as the input for the ARTn-MTP refined configurations.
As we first test the capability of ARTn coupled with MLIP to model amorphous systems, it is essential to explore several scenarios. A key consideration is the starting input configuration (see Table 1). Starting from a completely random state presents a challenge, particularly for larger systems. Ideally, the process should begin from a local minimum, which can be obtained through methods such as MD runs, CG relaxation, or ARTn coupled with a semi-empirical model. In this section, we present preliminary results that demonstrate the capability of ARTn-MTP to push the limits of relaxation for a configuration already optimized by ARTn coupled with the Stillinger-Weber (SW) potential 79.
A previously published a-Si model 22, 38 using ARTn with the SW potential 79, modified by Vink 80, served as input for an ARTn-MTP process. The 1000-atom ARTn-SW a-Si configuration (1000-aSi-SW initial) was refined through ARTn-MTP, resulting in the (1000-aSi-SW final) configuration. The goal was to assess the ARTn-MTP process’s ability to refine configurations optimized with a semi-empirical potential. The energy dropped by -23.25 meV/atom, and structural properties, shown in Fig. 9 and summarized in Tab. 3, reveal significant improvements. The number of dangling and floating bonds decreased, reducing total defects by 3%. Bond angles (Fig. 9 b) are closer to the ideal 109.5°, with a narrower spread in the final model. Bond lengths were concentrated around 2.38 Å, aligning with experimental values, and the radial distribution function in Fig. 9 c confirmed this. The 1000-aSi-SW final and 100k-GAP18 models showed better agreement with experimental data than the initial configuration. The structure factor, in Fig. 9 d, also matched experimental S(q). The excess energy ($\Delta E$) decreased from 0.2021 eV/atom (initial) to 0.1789 eV/atom (final), falling within the experimental range of 0.135 – 0.205 eV/atom 81. The final configuration, with lower excess energy, is more stable and realistic. The bulk modulus increased from 65.8573 GPa (initial) to 68.8690 GPa (final), indicating enhanced stiffness. These results demonstrate that the ARTn-MTP coupling scheme effectively relaxes a semi-empirically optimized amorphous configuration.
Properties 1000-aSi-SW initial 1000-aSi-SW final
Density (g/cm3) 2.20 2.24
Fourfold Si (%) 94.60 97.60
Threefold Si (%) 3.00 0.90
Fivefold Si (%) 2.40 1.50
Average bond lenght ($\sigma$) 2.38 (0.05) 2.38 (0.03)
Average bond angle ($\sigma$) 109.11 (10.91) 109.12 (11.26)
Average coordination number 3.99 4.01
$\Delta E$ (eV / atom) 0.20 0.18
Bulk modulus (GPa) 65.86 68.87
Crystallinity (%) 0.00 0.00
Tab. 3: Structural and eleastic properties for 1000-aSi-SW initial and 1000-aSi-SW final configurations: coordination defects, coordination number, average bond length, average bond angle, bulk modulus and energy difference per atom between amorphous state and crystalline Si. Standard deviations are provided in parentheses. The ART-SW optimized configurations (1000-aSi-SW initial) served as the input for the ARTn-MTP refined configurations (1000-aSi-SW final).
III.4 Stiffness
Fig. 10: Bulk modulus for ARTn-MTP generated models as well as models from the literature.
We calculated the bulk modulus $B$ of our models and of the models found in the literature. The results are presented in Fig. 10. We include the bulk modulus value obtained from density functional theory (DFT) simulations reported in Ref. 82, which is 82.5 GPa. For our 216-R2 configuration with 100% 4-fold coordination, we obtained a bulk modulus of 80.01 GPa. The 512-R2 model yields a bulk modulus of 77.53 GPa, closely matching the 512-WWW model’s value of 77.41 GPa. In contrast, other models such as 512-MD, 512-INV, and 512-RMC exhibit lower bulk modulus values of 71.18 GPa, 69.00 GPa, and 70.90 GPa, respectively. The 1000-R model demonstrates a bulk modulus of 76.46 GPa. Our largest models, the 4096-R-MD and 100k-GAP18, yield bulk modulus values of 77.93 GPa and 79.56 GPa, respectively. All of these values were obtained using our unified machine learning potential developed for the Si, O, and SiO2 systems. Other potentials may yield different results. To our knowledge, there are no clear experimental benchmark values for the bulk modulus of amorphous silicon.
III.5 Potential energy
Fig. 11: Excess energy between ART-MTP configurations (a-Si) and crystalline silicon (c-Si) as a function of the number of accepted events. A total of 4 events per atom was required to achieve 100% 4-fold coordination in 216-R2 a-Si model.
Models $\rho_{i}$ (g/cm3) $\Delta E$ (eV / atom) $\rho_{f}$ (g/cm3)
216-R1 2.23 0.16 2.24 (2.23)
216-R2 2.28 0.15 2.25 (2.24)
512-R1 2.23 0.17 2.23 (2.22)
512-R2 2.28 0.17 2.25 (2.24)
1000-R 2.28 0.19 2.25 (2.24)
4096-R-MD 2.28 0.16 2.26 (2.25)
Exp 2.2800 83 0.135 – 0.205 81 -
Tab. 4: Initial ($\rho_{i}$) and final ($\rho_{f}$) density for the ARTn-MTP generated a-Si model. Each model was randomly filled and the simulation box sizes was adjusted to match the initial density, which was maintained from the beginning to the end of the ARTn simulation. The final density was obtained by relaxing each of the generated models to 0.5 K as described in main test. The values in parentheses were obtained by relaxing each model at 300 K and 1 atm. The excess energy in eV per atom ($\Delta E$) relative to crystalline silicon is presented alongside the experimental excess energy.
Fig. 11 shows the relationship between energy decrease and the number of events for the model 216-R2. As can be seen, the energy decreases as the number of events increase. A detailed analysis of the events and key parameters is provided in the Nature of the Relaxation section. The experimental excess energy interval is indicated by the light red shaded region in the figure for reference. The inset of Fig. 11 shows the excess energy per atom for each of our models. These values are also presented in table 4. All models fall within the experimental excess energy interval 81, as detailed in the inset. Notably, the 216-R2 model, which initially contained over 70% coordination defects (see Fig. 2), was efficiently relaxed to achieve 100% 4-fold coordination and the lowest excess energy of 0.17 eV/atom with just 4 events per atom. To our knowledge, with the exception of a few Keating-potential-based WWW models 10, 78, 23, no other method, including DFT 84, has yielded a model of a-Si with no coordination defects.
III.6 Density
We computed the density of our models at 0 K. The final density was obtained by running a 50 ps molecular dynamics simulation, cooling the system from 300 K to 0.5 K, and then holding it at 0.5 K for an additional 50 ps, while barostatting the system at 0 Pa. In separate simulations, we also equilibrated each of the aforementioned models at 300 K and 1 atm for 100 ps. The results are presented in table 4. The ART-MTP a-Si density values at 0 K are 0.71% to 1.94% larger than the experimental density measured at room temperature. ART-MTP a-Si density at 300 K and 1 atm is characterized by density values 1.18% to 2.43% larger than experimental densities. The MTP potential is trained using a generalized gradient approximation exchange-correlation functional, which is known to overestimate the density of crystalline silicon by 2 % 85. Retraining the MTP on a higher-accuracy exchange-correlation functional, such as the restored regularized strongly constrained and appropriately normed meta-generalized gradient approximation 86 is likely to solve this small discrepancy with experimental values.
III.7 Nature of the relaxation
Fig. 12: Analysis of accepted Events in ARTn-MTP optimization: a) Barrier distribution, b) square root of squared atomic displacement versus energy barrier, c) energy barrier as a function of energy decrease, and d) energy barrier versus active atoms during events
The primary goal is not only to generate high-quality amorphous structures but also to understand the underlying physics of the relaxation process in amorphous systems over time. Thus, to understand the relaxation process in our amorphous system, we analyze simulation data, focusing on key parameters such as barrier distribution, atomic displacements, active atoms, and energy decrease, considering only accepted events as detailed in Fig. 12. These parameters can help provide insights into the nature and mechanism of the relaxation process. The data were collected under normal deformation, excluding shear deformation during the simulation. We first examine the correlation between atomic displacement and energy barriers. A clear correlation is observed: larger displacements are typically associated with higher energy barriers and more active atoms, as shown in Fig. 12 (a). To further validate this, we calculate the correlation between the square of the atomic displacement and the energy barriers. According to Hooke’s law, we expect a quadratic relationship between energy and displacement. When we plot energy against the square of the displacement, we expect a linear relationship, which our results confirm, yielding a strong correlation. Next, we find no significant correlation between the energy barriers and the energy gain, nor between energy gain and the number of active atoms, which is consistent with previous studies 87. Furthermore, no correlation exists between the number of atoms displaced up to the saddle point and the final state. As illustrated in Fig. 12(d), a correlation exists between energy barriers and the number of active atoms, with higher energy barriers typically involving a greater number of active atoms. On average, the number of active atoms varies from 10 to 30, with a mean of around 20 atoms that moves by more than 0.1 Å per event. This suggests that each event is localized, typically involving a cutoff radius of less than 7 Å. The microstructural changes primarily occur within a few neighboring cells, indicating a localized relaxation process. The relationship between the aforementioned parameters and the number of randomly displaced atoms at the start of ARTn searches (which depends on the cutoff and the central atom) was not investigated in this works.
III.8 Computational cost and nonlinear optimization process
Fig. 13: Nonlinear optimization effects characterized by excess energy per atom and number of events computed from independent runs with identical input Configuration
One of the primary challenges in computational materials modeling is finding the right balance between accuracy and computational cost. In this work, we present the first-ever coupling between ARTn and machine learning potentials, and we document the computational cost involved in preparing our amorphous system.
For the system of 216 atoms, the optimization of configurations required 0.0615 core-years and involved 55.20 millions force evaluations. Notably, one of the 216 atoms was optimized to achieve 100% 4-fold coordination. For the larger systems, the computational costs were as follows: the 512-atom system took 0.22 core-years, and the 1000-atom system required 0.34 core-years. Finally, for the 4096-atom system, we recorded 0.09 core-years to refine the MD configuration, with 37.25 million force evaluations. As mentioned in previous sections, this configuration achieved 98.05% 4-fold coordination and an excess energy of 0.15 eV per atom, which is closer to the experimentally observed low excess energy value.
The nonlinear optimization process within the ARTn framework was explored using a randomly generated 216-atom input configuration, with independent runs conducted using the same number of steps. Data were collected from these independent trajectories. Fig. 13 illustrates the relationship between excess energy levels and the number of accepted events for each trajectory. The results reveal varying numbers of accepted events and relaxation levels across the trajectories. Among the independent runs, the lowest number of accepted events was 45, while the highest was 113. Despite this difference, both systems achieved nearly identical excess energy levels. The mean number of accepted events was approximately 75, with most systems exhibiting higher excess energy than the system with 92 accepted events, which showed the lowest excess energy. These findings are indicative of a nonlinear optimization process.
IV Conclusions
In this paper, we investigated a-Si using ARTn coupled with a MTP. Seven models of varying sizes were constructed. They displayed excellent agreement with both experimental data and data obtained using a state-of-the-art 100,000-atom MD-based model leveraging a different MLIP. The proposed ARTn-MTP optimizer can further refine and relax configurations that have already been previously optimized using ARTn with a semi-empirical model. Additionally, the relaxation level of a-Si generated through standard molecular dynamics simulations via the melt-quench process can also be improved using the ARTn-MTP coupling scheme. We have generated high-quality amorphous structures of various sizes, including 216, 512, and 1000 atoms, without any traces of local crystallinity. We also found that starting the ARTn simulation with a randomly filled configuration results in an optimized structure with less crystallinity compared to initiating the simulation with an MD prerelaxed configuration. The question of local crystallinity, its link with defects and the ideal amorphous structure has not been addressed previously, due to the lack of good quality models generated through different pathways (non-dynamic, such as with ARTn, and melt-and-quench, for example). The results presented here raise numerous related research questions that could further enhance our understanding of a-Si and its fabrication. Future research could focus on increasing system sizes while minimizing coordination and ring defects, for example, by targeting over-coordinated, under-coordinated, and high-energy atoms during event searches. Further studies will investigate the origins of crystallinity, explore its formation during the preparation of amorphous silicon, and examine methods for producing large-scale amorphous a-Si free of crystallinity.
It is now clear that ARTn coupled with MLIPs is a very effective approach to model a-Si. However, if the primary goal is to replicate experimental conditions as accurately as possible, further innovation is necessary. Experimentally, a-Si is produced mainly by chemical vapor deposition and ion implantation; representative elementary volumes range from scales of tens of nanometers to micrometers, and processing times on the order of seconds to hours. Accessing such time- and length-scales would enable both applied studies–e.g. device-scale simulations–and fundamental studies–e.g. assessing the degree of hyperuniformity of a-Si. Although modeling large systems is straightforward thanks to parallel computing, reaching long timescales remains an elusive challenge. While there are examples of ARTn-based techniques reaching such timescales in irradiated c-Si and a-Si 88, 89, 90, 91, 92, further development is necessary in order for these methods to routinely achieve such outcome.
Acknowledgements
We thank the Digital Research Alliance of Canada for generous allocation of compute resources. Financial support was provided by the Natural Sciences and Engineering Research Council of Canada (NSERC) and the Association canadienne-française pour l’avancement des sciences (ACFAS), and the Canada Research Chair on Sustainable Multifunctional Construction Materials.
Author contributions
L.K.B. initiated and coordinated the research project. K.Z. performed all ARTn-MTP and molecular dynamics simulations, as well as the analyses presented in this work. All authors—K.Z., H.S., N.M., L.K.B., and C.O.P.—contributed to the writing of the paper. C.O.P., N.M., and L.K.B. supervised K.Z., while C.O.P. and L.K.B. provided the necessary computing resources.
Competing interests
The authors declare no conflict of interest.

### v2

@@
 Thermal annealing 18, 20, 13, 21 often refines these imperfections, enhancing material quality.
-Extensive theoretical research has focused on developing realistic structural models for a-Si 10, 22, 23, 24, 25, 26, 27.
+Additionally, a recent study 22 has raised an important issue that questions the definition of the true model of a-Si.
+This study explored the footprints of paracrystallinity in amorphous silicon using molecular dynamics simulations with a machine learning potential 22, 23 similar to the one employed herein.
+As reported by the authors, while optimal models of a-Si are traditionally based on the continuous random network model, some groups have proposed that real a-Si could be better described as ’paracrystalline’.
+However, since both optimal types of structure conflict with experimental observations 24, 25, 26, 27, the authors have further investigated the paracrystalline model, concluding that realistic experimental models of a-Si may contain some degree of crystalline grain traces.
+The observed degree of crystallinity is likely influenced by the preparation method—i.e.
+MD.
+Notably, while the degree of crystallinity was examined, its relationship to coordination defects was not.
+In other words, the central question remains how to balance crystallinity and coordination defects when preparing a-Si.
+For instance, a high cooling rate in melt-quench simulations will likely lead to a higher number of coordination defects and lower crystallinity, while a low cooling rate will most likely yield fewer coordination defects and higher crystalline-like environments.
+Extensive theoretical research has focused on developing realistic structural models for a-Si 10, 28, 29, 30, 31, 32, 33.
 Various methods for simulating materials, including a-Si, can be categorized into dynamic and non-dynamic approaches 7, 8.
-Dynamic methods, traditionally based on density functional theory (DFT) 28 or semiempirical force fields 29, primarily use melt-quench processes.
-However, these methods face challenges: experimental a-Si cannot be produced by melt-and-quench, DFT simulations are limited in system size and quenching rate, and semi-empirical force fields are often compromised by limited accuracy and transferability issues.
-Non-dynamic methods include the Reverse Monte Carlo (RMC) approach 30, 31, 32, 33, 34.
+Dynamic methods, traditionally based on density functional theory (DFT) 34 or semiempirical force fields 35, primarily use melt-quench processes.
+However, these methods face challenges: DFT simulations are limited in system size and quenching rate, and semi-empirical force fields are often compromised by limited accuracy and transferability issues.
+Non-dynamic methods include the Reverse Monte Carlo (RMC) approach 36, 37, 38, 39, 40.
 In RMC, the computer-generated a-Si structure is randomly modified in order for its simulated structure factor and radial distribution function g(r) to match experimental data.
@@
 To overcome this, hybrid RMC methods incorporate force field interactions, ensuring generated configurations are both consistent with experimental data and physically realistic.
-Variants like the force-enhanced atomic refinement 35, 36 and experimentally constrained molecular relaxation 37 utilize semi-empirical and DFT interactions, respectively.
-Other key non-dynamic methods include the Wooten-Winer-Weaire (WWW) bond-switching algorithm 10, which optimizes atomic configurations by minimizing energy, and the Activation-Relaxation Technique nouveau (ARTn) 22, 38, 39, 40, which combines activation and relaxation steps to discover low-energy structures.
-ARTn can be paired with various force fields and ab initio methods, making it suitable for applications that include relaxation of disordered systems and diffusion processes 41, 15, 42, 43, 44, 45, 46, 47, 48, 49.
+Variants like the force-enhanced atomic refinement 41, 42 and experimentally constrained molecular relaxation 43 utilize semi-empirical and DFT interactions, respectively.
+Other key non-dynamic methods include the Wooten-Winer-Weaire (WWW) bond-switching algorithm 10, which optimizes atomic configurations by minimizing energy, and the Activation-Relaxation Technique nouveau (ARTn) 28, 44, 45, 46, which combines activation and relaxation steps to discover low-energy structures.
+ARTn can be paired with various force fields and ab initio methods, making it suitable for applications that include relaxation of disordered systems and diffusion processes 47, 15, 48, 49, 50, 51, 52, 53, 54, 55.
 One of the key components in both dynamical and non-dynamical methods is the interaction potential.
-The ongoing challenge of balancing accuracy and computational cost in these methods has led to the development of machine learning interatomic potentials (MLIPs) 50, 51, 52.
+The ongoing challenge of balancing accuracy and computational cost in these methods has led to the development of machine learning interatomic potentials (MLIPs) 56, 57, 58.
 MLIPs combine descriptors of the atomic environment, regression methods, and DFT data.
-Many MLIPs frameworks have been developed, including artificial neural networks 50, 53, kernel-based methods 54, 55, and linear regression approaches 56, 57, among others.
-For a detailed description of MLIPs, including their implementation and testing, the reader is referred to the existing literature 52.
-In the case of a-Si, structures have been modeled using artificial neural networks 58, 59, 60 and Gaussian Approximation Potentials (GAP) 54, 55.
-In initial studies, the GAP potential was applied to a-Si systems containing between 512 and 4,096 atoms 24.
-Later, the GAP model was extended to larger systems of up to 100,000 atoms to investigate the the atomistic mechanisms underlying various structural transitions in disordered silicon 61.
-Other MLIPs, such as the Moment Tensor Potential 57 (MTP), have also been applied to a-Si 62.
+Many MLIPs frameworks have been developed, including artificial neural networks 56, 59, kernel-based methods 60, 61, and linear regression approaches 62, 63, among others.
+For a detailed description of MLIPs, including their implementation and testing, the reader is referred to the existing literature 58.
+In the case of a-Si, structures have been modeled using artificial neural networks 64, 65, 66 and Gaussian Approximation Potentials (GAP) 60, 61.
+In initial studies, the GAP potential was applied to a-Si systems containing between 512 and 4,096 atoms 30.
+Later, the GAP model was extended to larger systems of up to 100,000 atoms to investigate the the atomistic mechanisms underlying various structural transitions in disordered silicon 67.
+Other MLIPs, such as the Moment Tensor Potential 63 (MTP), have also been applied to a-Si 68.
 In these studies, melt-quench MD simulations were employed to generate a-Si structures.
-A natural next step is to couple ARTn 22, 38, 39, 40 with machine learning potential to generate a-Si models.
+A natural next step is to couple ARTn 28, 44, 45, 46 with machine learning potential to generate a-Si models.
 The objective of this paper is to generate a-Si models using ARTn-ML, which could enhance our understanding of the a-Si model and open new avenues for further research.
@@
 II.1 Activation Relaxation Technique nouveau
-ARTn 22, 38 is a computational method used in the study of activated processes, particularly in the context of diffusion and transitions between states in condensed matter systems.
+ARTn 28, 44 is a computational method used in the study of activated processes, particularly in the context of diffusion and transitions between states in condensed matter systems.
 ARTn is designed to simulate rare events that signify substantial changes in the microscopic structure of atomic systems, based on the efficient exploration of potential energy landscapes.
@@
 The concept is based on the understanding that within the configurational space, there is a distinct set of points representing saddle points and adjacent minima in the potential energy surface of the system.
-As the name implies, ARTn can be summarized as a two-step process 22: activation and relaxation.
+As the name implies, ARTn can be summarized as a two-step process 28: activation and relaxation.
 In the activation step, starting from a local energy minimum, the configuration is pushed toward a nearby first-order saddle point.
 In the relaxation step, the system is driven over the saddle point and then relaxed to a new energy minimum.
-The new configuration is accepted based on the Metropolis acceptance probability, which is given by $P_{\text{accept}}=\min\left(1,\exp\left(-\frac{\Delta E}{kT}\right)\right)$, where T is the fictitious temperature, set to 0.25 eV for our simulation, k is the Boltzmann constant, and $\Delta E$ is the energy difference between the old and new configurations 22, 63, 23.
+The new configuration is accepted based on the Metropolis acceptance probability, which is given by $P_{\text{accept}}=\min\left(1,\exp\left(-\frac{\Delta E}{kT}\right)\right)$, where T is the fictitious temperature, set to 0.25 eV for our simulation, k is the Boltzmann constant, and $\Delta E$ is the energy difference between the old and new configurations 28, 69, 29.
 The search is restricted to paths that include only saddle points.
 The ARTn method has demonstrated its value in revealing rare events that conventional simulation methods may struggle to capture due to the slow dynamics and relaxation inherent in these systems.
-Detailed descriptions of the method can be found elsewhere 22, 38, 39, 64, 65.
+Detailed descriptions of the method can be found elsewhere 28, 44, 45, 70, 71.
 II.2 A moment tensor potential for silicon
@@
 The simulation’s accuracy is directly influenced by the selection of the force field.
-Previous studies have relied on semi-empirical force field 42, 22, 66, 15, 67 and ab initio methods 68, 69, 44.
+Previous studies have relied on semi-empirical force field 48, 28, 72, 15, 73 and ab initio methods 74, 75, 50.
 Here we employ a machine learning-based forcefield.
-This potential was developed to jointly describe Si, O, and SiO2 systems 62, and is an improvement over a previously reported Si MTP 70.
-This potential is based on the MTP framework 57, 71.
-A detailed description of the development and implementation of this MTP can be found in our previous work 62, 70.
+This potential was developed to jointly describe Si, O, and SiO2 systems 68, and is an improvement over a previously reported Si MTP 76.
+This potential is based on the MTP framework 63, 77.
+The MTP potential was trained using data obtained from DFT calculations with the generalized gradient approximation (GGA-PBE) as the exchange-correlation functional 78.
+The MTP training set did not include amorphous silicon (a-Si) configurations.
+As explained in Ref.
+68, the MTP was validated against ab initio MD by performing a rapid melt-and-quench.
+Fast cooling (–8 ps) was followed by equilibration (10 ps) using either ab initio MD or MTP-MD.
+Both ab initio MD and MTP-MD led to a-Si models with very similar structural properties while repeating the same protocol with semi-empirical potentials led to substantially different results.
+The simulations presented herein can be seen as further validation and testing of this MTP.
+A detailed description of the development and implementation of this MTP can be found in our previous work 68, 76.
+II.3 Nomenclature of the models
+To name our models, including those from the literature, we used several techniques.
+Some models were named based on their creation procedure, while others were named according to the simulation method and the year the model was generated.
+For our models, we used the following naming conventions.
+Five of our models were named 216-R1, 216-R2, 512-R1, 512-R2, and 1000-R.
+Here, R stands for randomly generated initial structure, meaning the simulation box was randomly filled with atoms.
+When there are multiple models with the same number of atoms, the first model is named R1, and the second is named R2.
+If there is only one model for a given number of atoms, it is simply named R, as in 1000-R.
+One of our models was initially randomly filled and then pre-relaxed with MD simulation, so it was named 4096-R-MD.
+The models named 216-MD1, 216-MD2, 512-MD1, 512-MD2, 1000-MD, and 4096-MD refer to configurations that were entirely simulated using Molecular Dynamics (MD) with the ML (MTP) potential.
+The initial configurations are the same as those used for 216-R1, 216-R2, 512-R1, 512-R2, 1000-R, and 4096-R-MD, meaning the initial setups were randomly generated as described earlier.
+We found these input configurations to be a good starting point, as opposed to starting from a crystalline state and melting it.
+Models from the literature retain their original preparation methods, as indicated in the last column of Table I.
+For example, 216-FEAR indicates that the a-Si model was prepared using Force-enhanced atomic refinement, and 100k-GAP18 refers to a model with 100,000 atoms of a-Si, prepared using the Gaussian Approximation Potential (GAP) developed in 2018 for silicon materials.
 Models name Input configurations for a-Si preparation N atoms a-Si preparation methods
@@
 216-R2 Randomly filled 216 ARTn runs using MTP potential
-216-FEAR Randomly filled 216 Force hence atomic refinement 36
+216-MD1 Randomly filled 216 MD simulation using MTP potential
+216-MD2 Randomly filled 216 MD simulation using MTP potential
+216-FEAR Randomly filled 216 Force-enhanced atomic refinement 42
 512-R1 Randomly filled 512 ARTn runs using MTP potential
 512-R2 Randomly filled 512 ARTn runs using MTP potential
-512-MD Randomly filled 512 MD simulation using MTP potential
-512-RMC Randomly filled 512 Reverse monte carlo simulation 72, 24
-512-INV Randomly filled 512 Invariant environment refinement technique 72, 24
-512-FEAR Randomly filled 512 Force hence atomic refinement 36
-512-WWW Randomly filled 512 Bond switching algorithm 72, 24
+512-MD1 Randomly filled 512 MD simulation using MTP potential
+512-MD2 Randomly filled 512 MD simulation using MTP potential
+512-RMC Randomly filled 512 Reverse monte carlo simulation 79, 30
+512-INV Randomly filled 512 Invariant environment refinement technique 79, 30
+512-FEAR Randomly filled 512 Force-enhanced atomic refinement 42
+512-WWW Randomly filled 512 Bond switching algorithm 79, 30
 1000-R Randomly filled 1000 ARTn runs using MTP potential
+1000-MD Randomly filled 1000 MD simulation using MTP potential
 4096-R-MD Randomly filled + MD runs 4096 ARTn runs using MTP potential
+4096-MD Randomly filled + MD runs 4096 MD simulation using MTP potential
 4096-WWW Randomly filled 4096 Bond switching algorithm 42
-100k-GAP18 Unknown 100000 MD simulation using GAP potential 61
-1000-aSi-SW Amorphous configuration 22, 38, 68 1000 ARTn runs using MTP potential
+100k-GAP18 Unknown 100000 MD simulation using GAP potential 67
+1000-aSi-SW Amorphous configuration 28, 44, 74 1000 ARTn runs using MTP potential
 Tab.
@@
 1: Key structural elements characterizing a-Si: (a) bond angle ($\theta$), coordination defects such as dangling bonds (3-fold bonded atoms) and floating bonds (5-fold bonded atoms), and ring structures, including 4-, 5-, 6-, and 7-membered rings.
-(b) A 100% 4-fold fully coordinated amorphous structure comprising 216 atoms, generated by ARTn-MTP simulation , with 14.40 atoms in a crystalline environment.
+(b) A 100% 4-fold fully coordinated amorphous structure comprising 216 atoms, generated by ARTn-MTP simulation , with 14.4 % of atoms in a crystalline environment.
 The atoms shown in yellow within the box are in an amorphous environment, while those colored differently are in a crystalline environment, specifically cubic or hexagonal diamond.
-II.3 Protocol for input models preparation and simulation details
+II.4 Protocol for input models preparation and simulation details
 We prepared models containing 216, 512, 1000, and 4096 atoms.
@@
 The minimization is carried out using either the Steepest Descent (SD) or the Fast Inertial Relaxation Engine (FIRE) algorithm.
-In all ARTn-MTP simulations, we used a fictitious temperature of 0.25 eV in the Metropolis accept-reject criterion.
+In all ARTn-MTP simulations, we used a fictitious temperature of 0.25 eV in the Metropolis accept-reject criterion.The fictitious temperature is not a thermodynamic temperature of our system as we use ARTn here was a way to minimize total energy through activated events.
+Since no thermodynamical ensemble is therefore defined, this parameter serves as an artificial tool to aid in sampling configurations and achieving better convergence in the simulation.
 Each simulation started with event searches initiated with a local cut-off radius of 3 Å, which the algorithm used to identify the region around the selected central atom which will be deformed.
+The cut-off radius defines a spherical region around the central atom, within which both the selected central atom and its neighboring atoms are displaced during each event.
 This cut-off radius was progressively increased to a maximum cut-off 5.5 Å in order to expand the local zone of initial deformaton when the ARTn moves no longer led to energy decrease.
@@
 For defect levels at or below 5 %, only atoms with a non-four-fold coordination were allowed to serve as the central atom in ARTn moves.
-II.4 Structural analysis
-We employ a set of criteria to assess the quality of our amorphous structures.
-These criteria are derived from considerations of short-range structures, coordination defects, and intermediate-range order.
+For comparison purposes, we used the same initial (input) configurations as those for the ARTn-MTP simulations to run MD simulations with the MTP potential.
+For each MD model—216-MD1, 216-MD2, 512-MD1, 512-MD2, 1000-MD, and 4096-MD—the configuration was first heated to 2500 K over 30 ps, followed by equilibration at this temperature for 1 ns.
+The system was then cooled at a rate of 1 K/ps, and finally, the configuration was equilibrated at 300 K for 150 ps using the isothermal-isobaric ensemble (NPT).
+II.5 Structural analysis
+We use a set of criteria to assess the quality of our amorphous structures, based on short-range order, coordination defects, intermediate-range order, and the presence of crystalline grains, as illustrated in Fig.
+1 a).
 Understanding these different structural aspects is crucial in the study of materials properties, as they influence the physical, chemical, and mechanical behavior of materials.
-We conduct coordination analysis using OVITO 73 with a bond-length cutoff of 2.85 Å.
-The analysis of the ring size distribution, providing insights into intermediate-range order, is performed using the R.I.N.G.S code 74 based on the King criteria.
-We employed the ISAACS code 75 for the calculation of the structure factor while the radial distribution function and angular distribution function were determined using LAMMPS 76 through pseudo simulations.
-We also measure the amount of crystallinity in the models.
+We conduct coordination analysis using OVITO 80 with a bond-length cutoff of 2.85 Å.
+The analysis of the ring size distribution, providing insights into intermediate-range order, is performed using the R.I.N.G.S code 81 based on the King criteria.
+We employed the ISAACS code 82 for the calculation of the structure factor while the radial distribution function and angular distribution function were determined using LAMMPS 83 through pseudo simulations.
+We also quantify the crystallinity in the models, as indicated by the colored atoms (excluding the yellow atoms) in Fig.
+1 b.
 Here, crystallity refers to environments that are cubic or hexagonal diamond to second neighbour, all atoms at the center of this environment or within the first- and second-neighbour shell of a perfectly atom are considered as having crystallinity.
 We use method developed by Ref.
-77 as implemented in OVITO 73.
+84 as implemented in OVITO 80.
+To verify the crystalline content of our models, we used Common Neighbor Analysis (CNA) in OVITO, which offers four operation modes.
+Two are relevant for amorphous systems: the conventional CNA with a fixed cutoff and the Adaptive CNA with a variable cutoff.
+The conventional CNA is better suited for crystalline systems, while the Adaptive CNA is designed for multiphase systems, which is more relevant for amorphous systems with crystalline seeds.
+Both methods were applied to our models, but they are limited to lattice systems other than diamond and hexagonal diamond.
+Therefore, we then used the "Identify Diamond Structure" mode in OVITO, an extension of Adaptive CNA, which automatically determines the optimal cutoff radius for each particle.
+Finally, we varied the cutoff in conventional CNA but observed no changes in crystallinity.
 III Results and discussion
-Fig.
-2: Coordination numbers of input configurations and ARTn-MTP optimized configurations (final configurations).
-All non-4-fold coordinated atoms, including under- and over-coordinated atoms, are considered defective.
-The models analyzed include 216, 512 , 1000 and 4096 atoms.
 Tab.
@@
 216-R2 2.28 2.36 3.70 4.000 109.19 (10.02) 0.00 0.00 100.00 14.40
-216-FEAR 36 2.33 2.36 3.81 4.028 108.52 (15.59) 1.39 4.17 94.44
+216-MD1 2.25 2.37 3.83 4.018 109.08 (11.51) 0.00 1.85 98.15 0.00
+216-MD2 2.26 2.37 3.83 4.000 109.05 (10.60) 0.93 0.93 98.14 7.90
+216-FEAR 42 2.33 2.36 3.81 4.028 108.52 (15.59) 1.39 4.17 94.44
 512-R1 2.20 2.36 3.90 4.008 109.11 (11.39) 0.59 1.37 98.04 0.00
 512-R2 2.28 2.36 3.90 4.015 108.99 (11.76) 0.19 1.76 98.04 0.00
-512-MD 2.28 2.37 3.83 3.996 109.05 (10.95) 1.37 0.98 97.65 3.30
-512-RMC 72, 24 2.30 2.34 3.66 4.039 108.57 (14.16) 1.56 5.47 92.96 3.30
-512-INV 72, 24 2.31 2.34 3.86 4.054 108.56 (14.47) 1.95 7.42 90.62 0.00
-512-WWW 72, 24 2.27 2.34 3.94 4.008 109.05 (10.36) 0.19 0.98 98.82 10.00
-512-FEAR 36 2.33 2.35 3.82 4.008 1.17 2.73 95.90
-512-GAP18 24 2.27 2.36 3.76 4.004 109.20 (9.60) 0.58 0.98 98.44 6.4
-1000-R 2.28 2.37 3.76 4.020 108.94 (12.31) 0.6 2.6 97.20 0.00
+512-MD1 2.24 2.37 3.83 4.004 109.09 (11.10) 0.78 1.17 98.05 3.30
+512-MD2 2.24 2.37 3.90 3.996 109.13 (10.44) 1.37 0.98 97.65 3.30
+512-RMC 79, 30 2.30 2.34 3.66 4.039 108.57 (14.16) 1.56 5.47 92.96 3.30
+512-INV 79, 30 2.31 2.34 3.86 4.054 108.56 (14.47) 1.95 7.42 90.62 0.00
+512-WWW 79, 30 2.27 2.34 3.94 4.008 109.05 (10.36) 0.19 0.98 98.82 10.00
+512-FEAR 42 2.33 2.35 3.82 4.008 1.17 2.73 95.90
+512-GAP18 30 2.27 2.36 3.76 4.004 109.20 (9.60) 0.58 0.98 98.44 6.40
+1000-R 2.28 2.37 3.76 4.020 108.94 (12.31) 0.60 2.60 97.20 0.00
+1000-MD 2.25 2.37 3.83 4.016 109.98 (11.25) 0.70 2.30 97.00 13.30
 4096-R-MD 2.28 2.37 3.83 4.012 109.14 (10.65) 0.51 1.44 98.05 6.40
-4096-WWW 36 2.33 2.36 3.78 4.004 0.05 0.49 99.46
-100k-GAP18 61 2.27 2.36 3.83 4.003 109.17 (9.99) 0.69 0.95 98.35 6.5
+4096-MD 2.25 2.37 3.90 4.009 109.06 (11.04) 0.78 1.66 97.56 11.70
+4096-WWW 42 2.33 2.36 3.78 4.004 0.05 0.49 99.46
+100k-GAP18 67 2.27 2.36 3.83 4.003 109.17 (9.99) 0.69 0.95 98.35 6.5
 Fig.
-3: Comparison of the local structural properties of input configurations (top panels: a, b, and c) with the final configuration (100% coordinated) (bottom panels: d, e, and f) generated using the ARTn-MTP coupling scheme.
+2: Comparison of the local structural properties of input configurations (top panels: a, b, and c) with the final configuration (100% coordinated) (bottom panels: d, e, and f) generated using the ARTn-MTP coupling scheme.
 The structural properties analyzed include angular distribution, radial distribution, and bond length distribution for a system comprising 216 atoms.
-Experimental data and data from the 100k-GAP18 model 61 are included to assess the quality of our 100% 4-fold coordinated 216-atom a-Si model
+Experimental data and data from the 100k-GAP18 model 67 are included to assess the quality of our 100% 4-fold coordinated 216-atom a-Si model
 Models of size 216 atoms, 512 atoms, 1000 atoms and 4096 atoms were generated as detailed in table 1.
-We analyze these configurations with respect to short-range order, medium-range order, electronic properties as well as mechanical properties.
-Fig.
-2 displays the defects in both the input and final configurations.
-The input configurations exhibit a higher prevalence of coordination defects than the final configurations: approximately 70% for the 216-R1, 216-R2, and 512-R1 models, compared to around 30% for the 512-R2 and 1000-R models, as illustrated by the blue bars.
-The green bars represent the total number of defects present in the final configurations.
+We analyze these configurations with respect to short-range order, medium-range order, as well as mechanical properties.
+First, we would like to emphasize that, the input configurations exhibit a higher prevalence of coordination defects than the final configurations: approximately 70% for the 216-R1, 216-R2, and 512-R1 models, compared to around 30% for the 512-R2 and 1000-R models.
 In all ARTn-MTP a-Si models, the final percentage of defective atoms remained below 2%, with the exception of the 1000-R and 4096-R-MD models, which exhibit a final defect percentage of approximately 3%.
-The initial 4-fold coordination number (shown by the orange bars in Fig.
-2) is approximately 30% for the 216-R1, 216-R2, and 512-R1 models, and around 68% for the 512-R2 and 1000-R models.
-In contrast, the final configurations display markedly higher 4-fold coordination numbers (depicted by the red bars in Fig.
-2 and Table 2), with values of 99.07%, 100%, 98.04%, 98.04%, 97.20%, and 97.80% for the 216-R1, 216-R2, 512-R1, 512-R2, 1000-R and 4096-R-MD models, respectively.
+The initial 4-fold coordination number is approximately 30% for the 216-R1, 216-R2, and 512-R1 models, and around 68% for the 512-R2 and 1000-R models.
+In contrast, the final configurations display markedly higher 4-fold coordination numbers (Table 2), with values of 99.07%, 100%, 98.04%, 98.04%, 97.20%, and 98.05% for the 216-R1, 216-R2, 512-R1, 512-R2, 1000-R and 4096-R-MD models, respectively.
 No atoms with fewer than 3 neighbors or more than 5 neighbors were observed.
-To further contextualize the number of coordination defects in the ARTn-MTP a-Si models, we present a comparison in Fig.
-5, which includes data from the literature as well.
+To further contextualize the number of coordination defects in the ARTn-MTP a-Si models, we present a comparison in Table 2, which includes data from MD simulation and the literature as well.
 First, all the models presented show no atoms with fewer than 3 neighbors or more than 5 neighbors.
 Secondly, all of the ARTn-MTP generated models exhibit fewer floating and dangling bonds compared to all other methods, except the WWW model and 100k-GAP18 model.
-Notably, the 216-R2 model is absent from Fig.
-5 because it contains 100% four-fold coordination.
-To our knowledge, no other a-Si models exhibit 100% four-fold coordination, except for some WWW models generated with a Keating-type potential 78, 23
+Notably, the 216-R2 model exhibit 100% four-fold coordination.
+To our knowledge, no other a-Si models exhibit 100% four-fold coordination, except for some WWW models generated with a Keating-type potential 85, 29
 The last column of Table 2 presents the fraction of atoms in a crystalline environment (cubic or hexagonal diamond).
 As shown, most of the generated a-Si models, particularly those from MD simulations, contain some proportion of crystallinity.
-Although WWWW is known to produce large a-Si models without coordination defects, the WWW a-Si models still exhibit traces of crystallinity.
+Although WWW is known to produce large a-Si models without coordination defects, the WWW a-Si models still exhibit traces of crystallinity.
 We identified four models—216-R1, 512-R1, 512-R2, and 1000-R—that are free from crystallinity.
@@
 This can lead to understanding why, for example, amorphous silicon cannot be produced experimentally from the liquid phase and crystallizes well below melting as well as whether the CRN model corresponds really to ideal a-Si.
-The ML potential 62, 70, which leads to high-quality models, makes further study on this question now possible.
+The ML potential 68, 76, which leads to high-quality models, makes further study on this question now possible.
 Fig.
-4: Local structural properties of ARTn-MTP generated a-Si models: (a) radial distribution, (b) angular distribution, and (c) bond length distributions.
+3: Local structural properties of ARTn-MTP generated a-Si models: (a) radial distribution, (b) angular distribution, and (c) bond length distributions.
 Also shown are the properties obtained from experimental data and the 100k-GAP18 atom simulations.
 Fig.
-5: Proportion of coordination defects in various a-Si models generated by different methods.
-216-R1 has no floating bond and 216-R2 is 100 % fully 4-fold coordinated.
-No atoms with fewer than 3 neighbors or more than 5 neighbors have been recorded in any model.
-Fig. 6: FWHM of the bond-angle distribution. The FWHM values are computed by fitting a Gaussian distribution to the bond angle distributions. The red line denotes the experimental value of FWHM 20.
+4: FWHM of the bond-angle distribution.
+The FWHM values are computed by fitting a Gaussian distribution to the bond angle distributions.
+The bars are hatched for our purely MD-generated a-Si models.
+The red line denotes the experimental value of FWHM 20.
 III.1 Short-range order structural analysis of the models
@@
 2 illustrates the structural properties of the 216-R2 models, including the angular distribution function, radial distribution function, and bond length distribution.
-Panels a), b), and c) display these properties for the randomly packed input configuration (216-R2), while panels d), e), and f) show the properties of the final configurations (216-R2) produced by ARTn-MTP.
+Panels a, b, and c display these properties for the randomly packed input configuration (216-R2), while panels d, e, and f show the properties of the final configurations (216-R2) produced by ARTn-MTP.
 The final 216-R2 configuration is 100% fully coordinated, with no coordination defects present.
 As shown in Fig.
-3 a), b), and c), the data for the input configurations are randomly distributed and widespread.
+2 a, b, and c, the data for the input configurations are randomly distributed and widespread.
 In contrast, the data for the final configurations, depicted in Fig.
-3 d), e), and f), follow well-defined distributions.
+2 d, e, and f, follow well-defined distributions.
 The shape of the radial distribution function matches the experimental data, though the peak is somewhat overestimated.
@@
 The half width at half-maximum (HWHM) is 11.73°, which is similar to the experimental HWHM value of 11° 20.
-For the bond length, the data for the input configuration (3 c) are widespread.
-In contrast, for the final configuration (3 f), the bond lengths follow a Gaussian distribution centered around the experimental value of 2.35 Å.
+For the bond length, the data for the input configuration (Fig.
+2 c are widespread.
+In contrast, for the final configuration (Fig.
+2 f, the bond lengths follow a Gaussian distribution centered around the experimental value of 2.35 Å.
 The mean bond length is 2.36 Å, with a spread of 0.03 Å.
@@
 As shown in Fig.
-3 (d), (e), and (f), these properties align well with those of the larger 100k-GAP18 model.
+2 d, e, and f, these properties align well with those of the larger 100k-GAP18 model.
 This demonstrates the high quality of our fully coordinated a-Si model.
@@
 Since amorphous silicon features local tetrahedral bonding, we quantify the deviation from ideal tetrahedral geometry by calculating the FWHM for all models included in this study, as presented in Fig.
-6.
+4.
 The FWHM values for all our models closely align with the experimental data, as well as with the WWW model and the 100k-GAP18 models.
@@
 […]
-Fig. 7: Structure factor for ARTn-MTP generated models compared with that obtained from experimental data and 100k-GAP18 atom simulations.
-Fig. 8: Ring statistics for ARTn-MTP generated models as well as those derived from the 512-WWW model and 100k-GAP18 atom models.
+Fig. 5: Structure factor for ARTn-MTP generated models compared with that obtained from experimental data from 13 and 100k-GAP18 atom simulations.
+Fig.
+6: Ring statistics for ARTn-MTP generated models as well as those derived from the 512-WWW model and 100k-GAP18 atom models.
+The total ring count is normalized by the number of atoms in the model.
 Ring statistics are illustrated in Fig.
-8.
+6.
 Crystalline silicon, characterized by its diamond-type structure, primarily features 6-membered rings, while amorphous silicon tends to favor rings of sizes 5, 6, and 7.
@@
 As shown in Fig.
-8 (a-f), each model predominantly contains 5, 6, and 7-membered rings, consistent with existing literature.
+6 a-f, each model predominantly contains 5, 6, and 7-membered rings, consistent with existing literature.
 These ring sizes are energetically favorable for a-Si.
 Furthermore, all models exhibit a few larger ring defects with sizes greater than 7, as detailed in Fig.
-8.
+6.
 III.3 Refinement of ART-SW optimized model: 1000-aSi-SW
 Fig.
-9: Local structural properties of ARTn-MTP refined configurations (1000-aSi-SW final) compared to ART-SW optimized configurations (1000-aSi-SW initial).
+7: Local structural properties of ARTn-MTP refined configurations (1000-aSi-SW final) compared to ART-SW optimized configurations (1000-aSi-SW initial).
 a) Initial and final counts of 3-fold and 5-fold defects, b) bond angle distribution function, c) radial distribution function, and d) structure factor.
-For figures (b), (c), and (d), experimental data and results from the 100k-GAP-18 atom model are included to assess the quality of both the initial and final configurations of the 1000-aSi-SW model.
+For figures (b), (c), and (d), experimental data 13 and results from the 100k-GAP-18 atom model are included to assess the quality of both the initial and final configurations of the 1000-aSi-SW model.
 The ART-SW optimized configurations served as the input for the ARTn-MTP refined configurations.
@@
 Ideally, the process should begin from a local minimum, which can be obtained through methods such as MD runs, CG relaxation, or ARTn coupled with a semi-empirical model.
-In this section, we present preliminary results that demonstrate the capability of ARTn-MTP to push the limits of relaxation for a configuration already optimized by ARTn coupled with the Stillinger-Weber (SW) potential 79.
-A previously published a-Si model 22, 38 using ARTn with the SW potential 79, modified by Vink 80, served as input for an ARTn-MTP process.
+In this section, we present preliminary results that demonstrate the capability of ARTn-MTP to push the limits of relaxation for a configuration already optimized by ARTn coupled with the Stillinger-Weber (SW) potential 86.
+A previously published a-Si model 28, 44 using ARTn with the SW potential 86, modified by Vink 87, served as input for an ARTn-MTP process.
 The 1000-atom ARTn-SW a-Si configuration (1000-aSi-SW initial) was refined through ARTn-MTP, resulting in the (1000-aSi-SW final) configuration.
@@
 7 d, also matched experimental S(q).
-The excess energy ($\Delta E$) decreased from 0.2021 eV/atom (initial) to 0.1789 eV/atom (final), falling within the experimental range of 0.135 – 0.205 eV/atom 81.
+The excess energy ($\Delta E$) decreased from 0.2021 eV/atom (initial) to 0.1789 eV/atom (final), falling within the experimental range of 0.135 – 0.205 eV/atom 88.
 The final configuration, with lower excess energy, is more stable and realistic.
@@
 III.4 Stiffness
-Fig. 10: Bulk modulus for ARTn-MTP generated models as well as models from the literature.
+Fig.
+8: Mechanical properties of the models: a) Bulk modulus of ARTn-MTP generated models, MD generated models (hatched bars), and literature models.b) Elastic constants of ARTn-MTP generated a-Si models and MD simulated models (hatched bars).
 We calculated the bulk modulus $B$ of our models and of the models found in the literature.
 The results are presented in Fig.
-10.
+8 a.
 We include the bulk modulus value obtained from density functional theory (DFT) simulations reported in Ref.
-82, which is 82.5 GPa.
+89, which is 82.5 GPa.
+The DFT value was computed using the local density approximation (LDA) and hard, norm-conserving pseudopotentials 89.
 For our 216-R2 configuration with 100% 4-fold coordination, we obtained a bulk modulus of 80.01 GPa.
@@
 The 1000-R model demonstrates a bulk modulus of 76.46 GPa.
-Our largest models, the 4096-R-MD and 100k-GAP18, yield bulk modulus values of 77.93 GPa and 79.56 GPa, respectively.
-All of these values were obtained using our unified machine learning potential developed for the Si, O, and SiO2 systems.
+Our largest models, the 4096-R-MD and 100k-GAP18, yield bulk modulus values of 77.93 GPa and 79.56 GPa, respectively.For our a-Si models generated solely using MD simulation, the bulk moduli are as follows: 76.54 GPa for 216-MD1, 76.41 GPa for 216-MD2, 75.51 GPa for 512-MD1, 73.77 GPa for 512-MD2, 73.88 GPa for 1000-MD, and 74.66 GPa for 4096-MD.
+With the exception of 216-MD2, which shows no crystallinity, the other MD models contain significant proportions of crystallinity.
+The bulk modulus values presented here reflect the quality of our amorphous models.
+It is important to note that the bulk modulus for amorphous materials depends heavily on the preparation method and the inherent lack of periodicity in the structure.
+As a result, the value of the bulk modulus is not unique and varies across different models.
+The average bulk modulus of all the models, including our ARTn models, MD models, and those from the literature, is 75.47 GPa, with a spread of 3.63 GPa.
+All the bulk modulus values, including the mean, are notably lower than the bulk modulus of crystalline silicon, which is 98.3 GPa 90.
+A value of 59 GPa for the experimental bulk modulus of hydrogenated amorphous silicon (a-Si:H) has also been reported 91.
+This value is significantly lower than the bulk modulus values observed in all of the models presented.
+All these values were obtained using our unified machine learning potential developed for the Si, O, and SiO2 systems.
 Other potentials may yield different results.
 To our knowledge, there are no clear experimental benchmark values for the bulk modulus of amorphous silicon.
+In addition, we also assessed the elastic constants of all our models, and the results are presented in Fig.
+8 b.
+As can be noted, strong dispersion is not noticeable.
+For our ARTn-MPT models, the mean values (with spreads) are as follows: 118.46 GPa (4.77 GPa) for C11, 53.99 GPa (0.00 GPa) for C12, and 32.05 GPa (0.79 GPa) for C44.
+For our a-Si models generated solely with MD, the corresponding values are 119.87 GPa (1.45 GPa) for C11, 52.76 GPa (1.08 GPa) for C12, and 33.24 GPa (0.38 GPa) for C44.
+These values lie within the same range as those reported in previous works: 121 GPa for C11, 43 GPa for C12, 40 GPa for C44, and 69 GPa for Bulk modulus 30.
+Thus, while mechanical properties are highly dependent on the quality of the potential, we can argue that the computed values presented here fall within the range of the true mechanical properties of real a-Si models.
 III.5 Potential energy
 Fig.
-11: Excess energy between ART-MTP configurations (a-Si) and crystalline silicon (c-Si) as a function of the number of accepted events.
+9: a) Excess energy between ART-MTP configurations (a-Si) and crystalline silicon (c-Si) as a function of the number of accepted events.
 A total of 4 events per atom was required to achieve 100% 4-fold coordination in 216-R2 a-Si model.
+b) Excess energy for models generated using MD simulation solely
 Models $\rho_{i}$ (g/cm3) $\Delta E$ (eV / atom) $\rho_{f}$ (g/cm3)
@@
 216-R2 2.28 0.15 2.25 (2.24)
-512-R1 2.23 0.17 2.23 (2.22)
-512-R2 2.28 0.17 2.25 (2.24)
+512-R1 2.23 0.18 2.23 (2.22)
+512-R2 2.28 0.18 2.25 (2.24)
 1000-R 2.28 0.19 2.25 (2.24)
 4096-R-MD 2.28 0.16 2.26 (2.25)
-Exp 2.2800 83 0.135 – 0.205 81 -
+Exp 2.2800 92 0.135 – 0.205 88 -
 Tab.
@@
 Fig.
-11 shows the relationship between energy decrease and the number of events for the model 216-R2.
+9 a shows the relationship between energy decrease and the number of events for the model 216-R2.
 As can be seen, the energy decreases as the number of events increase.
@@
 The inset of Fig.
-11 shows the excess energy per atom for each of our models.
+9 a shows the excess energy per atom for each of our models.
 These values are also presented in table 4.
-All models fall within the experimental excess energy interval 81, as detailed in the inset.
-Notably, the 216-R2 model, which initially contained over 70% coordination defects (see Fig.
-2), was efficiently relaxed to achieve 100% 4-fold coordination and the lowest excess energy of 0.17 eV/atom with just 4 events per atom.
-To our knowledge, with the exception of a few Keating-potential-based WWW models 10, 78, 23, no other method, including DFT 84, has yielded a model of a-Si with no coordination defects.
+All models fall within the experimental excess energy interval 88, as detailed in the inset.
+Notably, the 216-R2 model, which initially contained over 70% coordination defects, was efficiently relaxed to achieve 100% 4-fold coordination and the lowest excess energy of 0.17 eV/atom with just 4 events per atom.
+To our knowledge, with the exception of a few Keating-potential-based WWW models 10, 85, 29, no other method, including DFT 93, has yielded a model of a-Si with no coordination defects.
+Fig.
+9 b shows the evolution of excess energy as a function of simulation time, with the inset corresponding to the excess energy of the generated a-Si models.
+Once again, the excess energy falls within the experimental range, although the values are higher than those for the ARTn-MTP-generated a-Si models.
+The mean excess energy for the ARTn-MTP models is 0.1667 eV per atom, with a spread of 0.0130 eV per atom, while for the MD-generated models, the mean excess energy is 0.1905 eV per atom, with a spread of 0.0035 eV per atom.
+This demonstrates that the ARTn-MTP-generated a-Si models are more stable than those generated by MD-MTP.
 III.6 Density
@@
 The results are presented in table 4.
-The ART-MTP a-Si density values at 0 K are 0.71% to 1.94% larger than the experimental density measured at room temperature.
-ART-MTP a-Si density at 300 K and 1 atm is characterized by density values 1.18% to 2.43% larger than experimental densities.
-The MTP potential is trained using a generalized gradient approximation exchange-correlation functional, which is known to overestimate the density of crystalline silicon by 2 % 85.
-Retraining the MTP on a higher-accuracy exchange-correlation functional, such as the restored regularized strongly constrained and appropriately normed meta-generalized gradient approximation 86 is likely to solve this small discrepancy with experimental values.
+Recall that the calculated density and experimental density of crystalline Si are 2.28 g/cm³ (variable cell relaxation 0 K) and 2.33 g/cm³ (300 K), respectively.
+The calculated density was obtained using an MTP potential that was trained on DFT data ( 0 K calculations).
+Therefore, a reasonable comparison between the two densities could not be made due to the difference in temperature conditions.
+We recorded a mean deviation of 1.46% from the calculated value and 3.57% from the experimental value of crystalline silicon.
+The MTP potential is trained using a generalized gradient approximation exchange-correlation functional, which is known to overestimate the density of crystalline silicon by 2 % 94.
+Retraining the MTP on a higher-accuracy exchange-correlation functional, such as the restored regularized strongly constrained and appropriately normed meta-generalized gradient approximation 95 is likely to solve this small discrepancy with experimental values.
 III.7 Nature of the relaxation
 Fig.
-12: Analysis of accepted Events in ARTn-MTP optimization: a) Barrier distribution, b) square root of squared atomic displacement versus energy barrier, c) energy barrier as a function of energy decrease, and d) energy barrier versus active atoms during events
+10: Analysis of accepted Events in ARTn-MTP optimization: a) Barrier distribution, b) square root of squared atomic displacement versus energy barrier, c) energy barrier as a function of energy decrease, and d) energy barrier versus active atoms during events
 The primary goal is not only to generate high-quality amorphous structures but also to understand the underlying physics of the relaxation process in amorphous systems over time.
 Thus, to understand the relaxation process in our amorphous system, we analyze simulation data, focusing on key parameters such as barrier distribution, atomic displacements, active atoms, and energy decrease, considering only accepted events as detailed in Fig.
-12.
+10.
 These parameters can help provide insights into the nature and mechanism of the relaxation process.
@@
 A clear correlation is observed: larger displacements are typically associated with higher energy barriers and more active atoms, as shown in Fig.
-12 (a).
+10 a.
 To further validate this, we calculate the correlation between the square of the atomic displacement and the energy barriers.
@@
 When we plot energy against the square of the displacement, we expect a linear relationship, which our results confirm, yielding a strong correlation.
-Next, we find no significant correlation between the energy barriers and the energy gain, nor between energy gain and the number of active atoms, which is consistent with previous studies 87.
+A fit was performed to identify potential trends linking atomic displacement to the system’s energy.
+For instance, Hooke’s law is often applied in similar contexts, where atomic displacement within a certain range is used to determine the spring constant.
+In this case, a log-log fit indicates a power law exponent of 2.4 with a standard error of ±0.7, considering only the ARTn events in the 4,096-atom system.
+An analysis over all the ARTn events generated in this study reveals a power law exponent of 1.90 with a standard error of ±0.04.
+These results suggest that Hooke’s Law likely applies.
+Next, we find no significant correlation between the energy barriers and the energy gain, nor between energy gain and the number of active atoms, which is consistent with previous studies 96.
 Furthermore, no correlation exists between the number of atoms displaced up to the saddle point and the final state.
 As illustrated in Fig.
-12(d), a correlation exists between energy barriers and the number of active atoms, with higher energy barriers typically involving a greater number of active atoms.
+10 d, a correlation exists between energy barriers and the number of active atoms, with higher energy barriers typically involving a greater number of active atoms.
 On average, the number of active atoms varies from 10 to 30, with a mean of around 20 atoms that moves by more than 0.1 Å per event.
@@
 IV Conclusions
-In this paper, we investigated a-Si using ARTn coupled with a MTP.
-Seven models of varying sizes were constructed.
-They displayed excellent agreement with both experimental data and data obtained using a state-of-the-art 100,000-atom MD-based model leveraging a different MLIP.
-The proposed ARTn-MTP optimizer can further refine and relax configurations that have already been previously optimized using ARTn with a semi-empirical model.
-Additionally, the relaxation level of a-Si generated through standard molecular dynamics simulations via the melt-quench process can also be improved using the ARTn-MTP coupling scheme.
-We have generated high-quality amorphous structures of various sizes, including 216, 512, and 1000 atoms, without any traces of local crystallinity.
-We also found that starting the ARTn simulation with a randomly filled configuration results in an optimized structure with less crystallinity compared to initiating the simulation with an MD prerelaxed configuration.
-The question of local crystallinity, its link with defects and the ideal amorphous structure has not been addressed previously, due to the lack of good quality models generated through different pathways (non-dynamic, such as with ARTn, and melt-and-quench, for example).
-The results presented here raise numerous related research questions that could further enhance our understanding of a-Si and its fabrication.
-Future research could focus on increasing system sizes while minimizing coordination and ring defects, for example, by targeting over-coordinated, under-coordinated, and high-energy atoms during event searches.
-Further studies will investigate the origins of crystallinity, explore its formation during the preparation of amorphous silicon, and examine methods for producing large-scale amorphous a-Si free of crystallinity.
-It is now clear that ARTn coupled with MLIPs is a very effective approach to model a-Si.
-However, if the primary goal is to replicate experimental conditions as accurately as possible, further innovation is necessary.
-Experimentally, a-Si is produced mainly by chemical vapor deposition and ion implantation; representative elementary volumes range from scales of tens of nanometers to micrometers, and processing times on the order of seconds to hours.
-Accessing such time- and length-scales would enable both applied studies–e.g.
-device-scale simulations–and fundamental studies–e.g.
-assessing the degree of hyperuniformity of a-Si.
-Although modeling large systems is straightforward thanks to parallel computing, reaching long timescales remains an elusive challenge.
-While there are examples of ARTn-based techniques reaching such timescales in irradiated c-Si and a-Si 88, 89, 90, 91, 92, further development is necessary in order for these methods to routinely achieve such outcome.
+This paper investigates a-Si using ARTn coupled with a MTP.
+Seven models of different sizes were developed, showing excellent agreement with both experimental data and a 100,000-atom MD-based model using a different MLIP.
+The ARTn-MTP optimizer refines configurations optimized by ARTn with a semi-empirical model.
+It also improves the relaxation of a-Si generated through the melt-quench process.
+High-quality amorphous structures with 216, 512, and 1000 atoms were produced, all free of local crystallinity.
+Starting ARTn simulations with a randomly filled configuration yields less crystallinity compared to using an MD-prerelaxed configuration.
+Overall, ARTn-MTP outperforms MD-MTP in modeling a-Si, as MD-generated a-Si often shows significant crystallinity, whereas ARTn-MTP produces a-Si with minimal crystallinity and low coordination defects.
+The relationship between local crystallinity, defects, and ideal amorphous structures remains underexplored due to a lack of high-quality models from different pathways like ARTn and melt-quench.
+Our results highlight new research directions for improving a-Si understanding and fabrication.
+Future work could focus on increasing system sizes while minimizing defects by targeting over- and under-coordinated atoms.
+Additional studies will investigate crystallinity’s origins, its formation in a-Si, and methods for producing large-scale amorphous a-Si without crystallinity.
+While ARTn combined with MLIPs is effective for modeling a-Si, further innovation is needed to replicate experimental conditions more accurately.
+a-Si is mainly produced by chemical vapor deposition and ion implantation, with representative volumes ranging from tens of nanometers to micrometers and processing times from seconds to hours.
+Achieving these time- and length-scales is essential for both applied and fundamental studies.
+Despite advancements in parallel computing, reaching long timescales remains a challenge.
+Data availability
+The potential file and the a-Si models are publicly available at https://gitlab.com/Kazongogit/MTPu/-/tree/main/ART/Si?ref_type=heads.
+The MTP source code can be found at https://mlip.skoltech.ru/, and the ARTn source code is freely available through a request to Normand Mousseau.
 Acknowledgements